How much energy is required to convert 12.00 g of ice at -5.00 °C to liquid water at 2.50 °C?



1
Expert's answer
2021-04-25T06:12:28-0400

Converting "12.0g" of ice to "0\u00b0C"

"=12.0\times 10=1200calories"

Converting to water at "2.5\u00b0C"

"=12.0g\times 2.5=30 calories"

Latent heat of vaporization of water "=537calories\/g"

"=537\times 12=6444 calories"

Total heat of fission"=(1200+30+6444)=7674 calories"